OpenStudy (waheguru):

cant get the right anwer why -2u-8=5u-1

OpenStudy (waheguru):

can u explain teh steps

OpenStudy (waheguru):

neeed hrlp

OpenStudy (tcp2315):

-8+1=5u+2u -7=7u -1=u

OpenStudy (anonymous):

use subtraction and addition to move all the terms with u's in them to one side and al the terms without us to the other

OpenStudy (waheguru):

but if uplug that in itz not even

OpenStudy (waheguru):

the textbook says the answer is -3???

OpenStudy (mimi_x3):